1. Solve the nonhomogeneous linear system of differential equations d / 8 3 1 / -10 25 u + dt ( 3 8 ) | 10 l -25
Answer #1

0 )u+ (-19)+(25)+ > first we find eigenvalues 1A-dIl=0 8-)(8-)-(3) (3)=0 64-82-84+1?q=0 x²16x +55=0 2411-54+55=0 XCA-11)-5(4-256 RR, 1-3 6 :19 >) 2,-*=0 ) 2132 put =1, then 2011 --Va= acce 5(-)+czet (5) Let up: (61) ++ (8) up = (a) plug it in systemwe have (891+36, = -25)3 (391+85,= 25)8 24 94+ ab, -75 244, +64 b, 200 –55b, = -275 bl=5 Sait 3b, = -25 :9,=-5, b,=5 89,+305)